# Ego-Vision World Model for Humanoid Contact Planning

### v1

@@ III-C Value-Guided Sampling MPC
    Partial Observability and Physical Non-idealities: Robotic systems suffer from partial observability, as the full state, such as contact forces, is not observable. is not directly measured, and is subject to sensor noise and action delays, both of which degrade value estimation.
@@
Consider a standard MDP with state $s_{t}$ and action $a_{t}$ for the analysis of the variance reduction of our proposed method. We denote the estimation of the true $Q$ functions as $\hat{Q}$. When the $Q$ is available, we can leverage the Bellman principle to obtain the optimal control input:
@@
However, in practice, only the estimation $\hat{Q}$ is available, which may have a large variance. The imperfection of $\hat{Q}$ may lead to performance degradation. To mitigate this issue, instead of directly optimizing on the estimated Q function, we consider an $N$-step surrogate optimization:
$\hat{\pi}(s_{t})=\arg\max_{\{a_{0:N-1}\}}$ $\hat{J}_{N}=\frac{1}{N}\sum_{t=0}^{N-1}\hat{Q}(s_{t},a_{t})$ (11)
$\mathrm{s.t.}$ $\quad s_{t+1}=f(s_{t},a_{t})$
@@
Consider the residual between $\hat{Q}$ and $Q$ as $\epsilon_{Q}$, and
$\hat{Q}(s_{t},a_{t})=Q(s_{t},a_{t})+\epsilon_{Q}(s_{t},a_{t}).$
For shorthand notation, we use the subscript $(t)$ to denote the quantity at time step $t$. We have the surrogate loss function:
$\frac{1}{N}\sum_{t=0}^{N-1}\hat{Q}_{t}=\frac{1}{N}\sum^{N-1}_{t=0}\left(Q_{t}+\epsilon_{t}\right).$ (12)
As we use Monte Carlo estimation to compute $\hat{Q}$ as an unbiased estimation, the mean of $\epsilon_{t}$ is zero, and thus the variance on the RHS can be obtained by
$\operatorname{Cov}[\hat{J}_{N}]=\operatorname{Var}[\frac{1}{N}\sum^{N-1}_{t=0}\epsilon_{t}]=\frac{1}{N^{2}}\sum_{(i,j)}\operatorname{Cov}[\epsilon_{i},\epsilon_{j}].$ (13)
We assume that the variance of $\epsilon_{t},\forall t,$ satisfies:
$0\leq V_{\min}\leq\operatorname{Var}[\epsilon_{t}]\leq V_{\max}$ (Bounded Variance)
@@
$|\operatorname{Cov}[\epsilon_{i},\epsilon_{j}]|\leq\rho\operatorname{Var}[\epsilon_{i}]\operatorname{Var}[\epsilon_{j}].$ (Bounded Correlation)
The upper bound variance is achieved if all steps are positively correlated:
$\operatorname{Cov}[\hat{J}_{N}]\leq\frac{N+\rho(N^{2}-N)}{N^{2}}V_{\max}=:V_{\operatorname{ub}}.$ (14)
@@
$\operatorname{Cov}[\hat{J}_{N}]\geq\max\left\{\frac{NV_{\min}-\rho(N^{2}-N)V_{\max}}{N^{2}},0\right\}=:V_{\operatorname{lb}}.$ (15)
Thus, we have limit when $\rho\rightarrow 0:V_{\operatorname{ub}}\rightarrow\frac{V_{\max}}{N},V_{\operatorname{lb}}\rightarrow\frac{V_{\min}}{N}.$ As we compute $\hat{Q}$ from an offline dataset generated by random action in each step, the correlation between $\hat{Q}_{i}$ from different time steps is weaker than by sampling using a state feedback policy. Thus, it is possible that the $\rho$ is small and thus $\hat{J}_{N}$ has a substantially lower variance than $\hat{Q}$. Although additional bias is introduced because the surrogate loss does not preserve the local optimum, if the variance of $\hat{Q}$ dominates, this strategy can significantly improve the performance.
Based on the above analysis, we apply the surrogate objective function $\hat{J}_{N}$ from  (11) using the latent variables $h_{t}$ and $\hat{z}$ as the representation of the robot states. The optimal sequence, denoted $A_{t}^{*}=\{a_{t}^{*},a_{t+1}^{*},...,a_{t+N-1}^{*}\}$, is the one that maximizes our surrogate objective $\hat{J}_{N}$:
@@
As shown in Fig. 3, the framework’s memory is maintained in two latent states: a dynamics latent state $h_{t}$ and a current observation latent state $z_{t}$. The $h_{t}$ is computed by the RNN from the previous latent ($h_{t-1},z_{t-1}$) and the last action ($a_{t-1}$). The $z_{t}$ is then generated in one of two ways: when an observation $o_{t}$ is available, $z_{t}$ is inferred using both the latent state $h_{t}$ and the observation $o_{t}$; for future predictions, it is generated from the latent state $h_{t}$ alone.
Our world model also predicts the probability of robot failure, $\hat{d}_{t}$, such as falling; if this probability exceeds a threshold of 0.9, all subsequent value estimates for that trajectory are set to zero.
We use the Cross-Entropy Method (CEM) to find optimal action sequence $A_{t}^{*}$. Once identified, only the first action $a_{t}^{*}$ is executed. This iterative re-planning allows the robot to continuously incorporate feedback from the environment, enabling it to react to disturbances and correct for model inaccuracies in real-time. From Table I, we select a planning horizon $N=4$ as our default, as it provides the best overall performance across all tasks.
@@ IV Experiments
Our experimental platform is the Unitree G1 humanoid robot, equipped with a RealSense D435i camera. All quantitative analyses, including ablation studies and baseline comparisons, are conducted in a controlled simulation environment. All comparisons are conducted using a consistent set of training epochs and hyperparameters. The mean and standard deviation are then computed from ten independent trials across three different random seeds. We subsequently validate our approach with real-time experiments on a physical robot. We designed three core tasks to evaluate the model’s ability to plan and execute diverse contact-rich behaviors, including exploit contact for stability and avoid contact for safety.
@@
We compare our method with these baseline methods: (1) PPO: implementation in [2] . (2) ARWM: replace our framework with auto-regressive prediction training like [29, 30]. (3) Rew-MPC: replace objective function Eq. (16) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}$ from PlaNet [24]. (4) TD-MPC: replace objective function Eq. (16) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}+\gamma^{N}\hat{Q}$ from TD-MPC [23].
@@ IV-A Advantages of the Use of Offline data
To compare single-task training performance, both our method and the PPO baseline leverage an identical low-level controller that tracks commands for end-effector position ($p_{ee}^{\top}$) and body height ($h_{body}$).
As shown in Fig. 5, in three contact-reward-dominant tasks, our method completes the task using only 0.5M data steps. In contrast, PPO requires a significantly larger amount of data, especially in simulations that necessitate visual rendering, where it consumes considerably more time. While PPO can quickly match our efficiency in tasks with simple visual features and a stationary robot, such as “Block the Ball,” our method achieves better performance on tasks with more complex visual representations and significant changes in the robot’s viewpoint. For instance, in the “Traverse the Arch” task, where the robot’s perspective changes dramatically between standing and squatting positions, our method substantially outperforms PPO.
@@
Figure 7: Visualization of the world model’s prediction and planning process during the “Block the Ball” task. (a) The evolution of the Q-value map for the hand’s position in the X-Z plane. As the task unfolds (from Step 30 to 80, with the ball appearing at Step 60), the model dynamically updates its value estimates, with the high-value region (yellow) indicating the optimal location to intercept the object. (b) An open-loop prediction of future observations. Given an initial image, the model first reconstructs it (Horizon 0) and then generates a sequence of future frames (Horizon 2-16) by decoding its predicted latent states, visualizing its anticipation of the ball’s trajectory. It is worth noting that while our planner uses a shorter 4-step horizon, this visualization is for demonstrating the long-horizon physical coherence of our learned model.
@@ IV-B Analysis of Key Design Choices
To demonstrate the necessity of our design, as shown in the Table I, we found that greedily maximizing the value (i.e., the horizon-1 case) is infeasible. This approach causes the robot to be myopic, favoring the maintenance of its default position and ignoring future contacts. We observe that longer horizons (e.g., N=6N{=}6) degrade performance, likely because bias dominates from longer-term prediction (see Sec. III-C), whereas N=4N{=}4 strikes a bias–variance sweet spot.
@@
As shown in the Table I, Rew-MPC, which uses reward as the objective function, yielded suboptimal results due to partial observability, noise, action lag, and sparse contact, which make rewards difficult to predict. And TD-MPC, which uses TD-target to evaluate trajectories, also costs unstable results. We attribute this to TD-error methods converging to deceptive solutions, where low TD error might mask highly inaccurate value estimates. As argued in [38], this phenomenon is caused by bias cancellation and the existence of infinitely many suboptimal solutions that satisfy the Bellman equation on an incomplete offline dataset.
@@ IV-C Multi-Task Planning with Unified World Model
To evaluate the multi-task capability of our model, we trained a single model on a combined dataset from all tasks and compared it to the specialized single-task models from Table I. As shown in Fig. 6, the multi-task model achieves improved performance on two of the three tasks, with a minor drop in the “Block the Ball” task, which we attribute to its smaller reward scale.
@@ IV-D Model Interpretation and Visualization on Prediction
Long-Horizon Prediction: To verify the model’s ability to make long-horizon predictions, we visualize its open-loop rollouts. As illustrated in Fig. 7(b), given a single initial observation, the model first reconstructs it (Horizon 0) and then generates a sequence of expected future observations (Horizon 2-16) by decoding its predicted latent states. This predicted sequence clearly visualizes the model’s anticipation of the ball’s trajectory from left to right. This demonstrates that the model has learned a coherent and physically plausible model of the world, even without auto-regressive training and constrained solely by the joint-embedding loss. The blurriness in the generated frames is reasonable, as our highly compressed 32-dimensional latent space forces the model to extract only the most salient physical information
Contact-Directed Planning: Figure 7(a) visualizes how the model leverages its predictions for planning by showing the evolution of the objective function value in Eq. (11) for the hand’s target position. Early in the task (e.g., Step 30), the value map consistently encourages the hand to stay near a natural, energy-efficient default position. However, as the ball approaches and the plan solidifies (e.g., Step 60-70), a high-value region (yellow) emerges and sharpens, decisively guiding the robot’s hand toward the optimal contact point. This dynamic evolution of the value map showcases the model’s contact-directed reasoning, effectively forming an interpretable plan to achieve its objective.
@@ IV-E Real-World Validation
We deployed our method for real-time experiments on Unitree G1 with 25 Hz real-time planning and evaluated a batch with 1024 action trajectories with a planning horizon of 4 steps at each timestep. The desired base velocity $v$ was controlled by a human operator.
Our real-world deployments, shown in Fig. 4, included both single-task and multi-task models, both of which proved capable of completing their assigned tasks. The policy also demonstrated the ability to generalize to out-of-distribution (OOD) scenarios, such as blocking a previously unseen box. These experiments validate that our method can achieve agile and robust vision-based control. Crucially, the learned policy exhibits reactive, context-dependent behavior rather than overfitting to a single action pattern. For instance, in the ’Support the Wall’ task, the robot only braces its hands against the wall when actively disturbed and returns to a neutral stance once balance is recovered. Please check our supplementary video for more details.
@@ V Conclusion
By integrating a scalable ego-centric visual world model with value-guided sampling-based model predictive control, we demonstrate that humanoid robots can efficiently and robustly learn agile, contact-rich behaviors from offline, demonstration-free data, advancing data-efficient, vision-based planning for real-world robotic interaction.
@@ ACKNOWLEDGMENT
This work was supported in part by NSF Grant CMMI-1944722 and by the Robotics and AI Institute. M. Ghaffari receives support by AFOSR MURI FA9550-23-1-0400. The authors would like to thank Jiaze Cai, Yen-Jen Wang for their help in experiments. We are also grateful to Bike Zhang, Fangchen Liu, Chaoyi Pan, Junfeng Long and Yiyang Shao for their valuable discussions. This work was done during H. Liu’s visit to UC Berkeley.

### v2

@@ III-D Value-Guided Sampling MPC
    Partial Observability and Physical Non-idealities: Robotic systems suffer from partial observability, as the full state, such as contact forces, is not directly measured and is subject to sensor noise and action delays, both of which degrade value estimation.
@@
Consider a standard MDP with state $s_{t}$ and action $a_{t}$ for the analysis of the variance reduction of our proposed method. When the perfect $Q$ is available, we can leverage the Bellman principle to obtain the optimal control policy:
@@
However, in practice, only the estimation $\hat{Q}$ is available, which may have a large variance. The imperfection of $\hat{Q}$ may lead to performance degradation. To mitigate this issue, we consider an $N$-step surrogate optimization:
$\hat{\pi}(s_{t})=\arg\max_{\{a_{t:t+N-1}\}}$ $\hat{J}_{N}=\frac{1}{N}\sum_{k=0}^{N-1}\hat{Q}(s_{t+k},a_{t+k})$ (11)
$\mathrm{s.t.}$ $\quad s_{t+k+1}=f_{\phi}(s_{t+k},a_{t+k})$
@@
Consider the residual $\epsilon_{t+k}$ at time $(t+k)$ as:
$\epsilon_{k}:=\hat{Q}(s_{t+k},a_{t+k})-Q(s_{t+k},a_{t+k}).$ (12)
For shorthand notation, we use the subscript $(k)$ to denote the quantity at time step $t+k$. Consider the surrogate objective:
$\frac{1}{N}\sum_{k=0}^{N-1}\hat{Q}_{k}=\frac{1}{N}\sum^{N-1}_{k=0}\left(Q_{k}+\epsilon_{k}\right).$ (13)
Let $J_{N}:=\frac{1}{N}\sum_{k=0}^{N-1}Q_{k}$. As we use Monte Carlo returns to form $Q_{target}$ and train $\hat{Q}$ via regression, we have:
$\mathbb{E}[(J_{N}-\hat{J}_{N})^{2}]$ $=\mathbb{E}[(\frac{1}{N}\sum^{N-1}_{k=0}\epsilon_{k})^{2}]$ (14)
$=\frac{1}{N^{2}}\sum_{(i,j)}\operatorname{Cov}[\epsilon_{i},\epsilon_{j}]+\bar{\epsilon}^{2}.$
with $\bar{\epsilon}:=\mathbb{E}[\epsilon_{k}]$ denoting the bias introduced by estimating $Q_{k}$ corresponding to the infinite horizon MDP using truncated trajectories. We assume that the variance of $\epsilon_{k},\forall k,$ satisfies:
$0\leq\sigma_{\min}\leq\sqrt{\operatorname{Var}[\epsilon_{k}]}\leq\sigma_{\max},$ (Bounded Variance)
@@
$|\operatorname{Cov}[\epsilon_{i},\epsilon_{j}]|\leq\rho\sqrt{\operatorname{Var}[\epsilon_{i}]\operatorname{Var}[\epsilon_{j}]}.$ (Bounded Correlation)
The upper bound on the variance term is achieved if all steps are positively correlated:
$\mathbb{E}[(J_{N}-\hat{J}_{N})^{2}]-\bar{\epsilon}^{2}$ $\leq\frac{N+\rho(N^{2}-N)}{N^{2}}\sigma^{2}_{\max}$ (15)
$=:V_{\operatorname{ub}}.$
@@
$\mathbb{E}[(J_{N}-\hat{J}_{N})^{2}]-\bar{\epsilon}^{2}$ $\geq\max\{\frac{N\sigma^{2}_{\min}-\rho(N^{2}-N)\sigma^{2}_{\max}}{N^{2}},0\}$ (16)
$=:V_{\operatorname{lb}}.$
Thus, we have the limit when $\rho\rightarrow 0:V_{\operatorname{ub}}\rightarrow\frac{\sigma^{2}_{\max}}{N},V_{\operatorname{lb}}\rightarrow\frac{\sigma^{2}_{\min}}{N}$. As we compute $\hat{Q}$ from an offline dataset generated by random actions in each step, the correlation between $\hat{Q}_{i}$ from different time steps is weaker than by sampling using a state feedback policy. Thus, it is possible that $\rho$ is small and $\hat{J}_{N}$ has a substantially lower variance than $\hat{Q}$. Although the surrogate objective does not preserve a local optimum, if the variance of $\hat{Q}$ dominates, this strategy can significantly improve the performance.
Based on the above analysis, we apply the surrogate objective function $\hat{J}_{N}$ from  (11) using the latent $h_{t}$ and $\hat{z}_{t}$ as the representation of the robot states. The optimal sequence, denoted $A_{t}^{*}=\{a_{t}^{*},a_{t+1}^{*},...,a_{t+N-1}^{*}\}$, is the one that maximizes our surrogate objective $\hat{J}_{N}$:
@@
As shown in Fig. 3, the framework’s memory is maintained in two latent states: a dynamics latent state $h_{t}$ and a current observation latent state $z_{t}$. The $h_{t}$ is computed by the RNN from the previous latent ($h_{t-1},z_{t-1}$) and the last action ($a_{t-1}$). The $z_{t}$ is then generated in one of two ways: when an observation $o_{t}$ is available, $z_{t}$ is inferred using both the latent state $h_{t}$ and the observation $o_{t}$; for future predictions, it is generated from the latent state $h_{t}$ alone. Our world model also predicts the probability of robot failure, $\hat{d}_{t}$, such as falling; if this probability exceeds a threshold of 0.9, all subsequent value estimates for that trajectory are set to zero.
We use the Cross-Entropy Method (CEM) to find the optimal action sequence $A_{t}^{*}$. Once identified, only the first action $a_{t}^{*}$ is executed. This iterative re-planning allows the robot to continuously incorporate feedback from the environment, enabling it to react to disturbances and correct for model inaccuracies in real-time. From Table I, we select a planning horizon $N=4$ as our default, as it provides the best overall performance across all tasks.
@@ IV Experiments
Our experimental platform is the Unitree G1 humanoid robot, equipped with a RealSense D435i camera. All quantitative analyses, including ablation studies and baseline comparisons, are conducted in a controlled simulation environment. All comparisons are conducted using a consistent set of training epochs and hyperparameters. The mean and standard deviation are then computed from ten independent trials across three different random seeds. We subsequently validate our approach with real-time experiments on a physical robot. We designed three core tasks to evaluate the model’s ability to plan and execute diverse contact-rich behaviors, including exploiting contact for stability and avoiding contact for safety.
@@
We compare our method with these baseline methods: (1) PPO: an implementation in [26]. (2) ARWM: replaces our framework with auto-regressive prediction training like [13, 14]. (3) Rew-MPC: replaces objective function Eq. (17) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}$ from PlaNet [9]. (4) TD-MPC: replaces objective function Eq. (17) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}+\gamma^{N}\hat{Q}$ from TD-MPC [11].
@@ IV-A Advantages of the Use of Offline Data
As shown in Fig. 5, in three contact-reward-dominant tasks, our method completes the tasks using only 0.5M data steps. In contrast, PPO requires a significantly larger amount of data, especially in simulations that necessitate visual rendering, where it consumes considerably more time. While PPO can quickly match our efficiency in tasks with simple visual features and a stationary robot, such as “Block the Ball,” our method achieves better performance on tasks with more complex visual representations and significant changes in the robot’s viewpoint. For instance, in the “Traverse the Arch” task, where the robot’s perspective changes dramatically between standing and squatting positions, our method substantially outperforms PPO.
@@
Figure 7: Visualization of the world model’s prediction and planning process during the “Block the Ball” task. (a) The evolution of the Q-value map for the hand’s position in the X-Z plane. As the task unfolds (from Step 30 to 80, with the ball appearing at Step 60), the model dynamically updates its value estimates, with the high-value region (yellow) indicating the optimal location to intercept the object. (b) An open-loop prediction of future observations. Given an initial image, the model first reconstructs it (Horizon 0) and then generates a sequence of future frames (Horizon 2-16) by decoding its predicted latent states, visualizing its anticipation of the ball’s trajectory. It is worth noting that while our planner uses a shorter 4-step horizon, this visualization is for demonstrating the long-horizon physical intuition.
@@ IV-B Analysis of Key Design Choices
To demonstrate the necessity of our design, as shown in the Table I, we found that greedily maximizing the value (i.e., the horizon-1 case) is infeasible. This approach causes the robot to be myopic, favoring the maintenance of its default position and ignoring future contacts. We observe that longer horizons (e.g., N=6N{=}6) degrade performance, likely because bias dominates from longer-term prediction (see Sec. III-D), whereas N=4N{=}4 strikes a bias–variance sweet spot.
@@
As shown in the Table I, Rew-MPC, which uses reward as the objective function, yielded suboptimal results due to partial observability, noise, action lag, and sparse contact, which make rewards difficult to predict. And TD-MPC, which uses TD-target to evaluate trajectories, also produces unstable results. We attribute this to TD-error methods converging to deceptive solutions, where low TD error might mask highly inaccurate value estimates. As argued in [6], this phenomenon is caused by bias cancellation and the existence of infinitely many suboptimal solutions that satisfy the Bellman equation on an incomplete offline dataset.
@@ IV-C Multi-Task Planning with Unified World Model
To evaluate the multi-task capability, we trained a single model on a combined dataset from all tasks and compared it to the single-task models from Table I. As shown in Fig. 6, the multi-task model achieves improved performance on two of the three tasks, with a minor drop in the “Block the Ball” task, which we attribute to its smaller reward scale.
@@ IV-D Model Interpretation and Visualization on Prediction
Physical Intuition: To assess whether the learned model captures task-relevant physical properties, rather than collapsing to trivial solutions, we apply long-horizon open-loop rollouts for verification. As shown in Fig. 7 (b), the rollout preserves the parabolic motion of the ball over 16 steps, indicating that the latent states retain physically meaningful structure. Importantly, exhibiting long-horizon physical intuition in open-loop rollouts does not imply that increasing the MPC planning horizon will necessarily improve control performance. In sampling-based MPC, longer horizons typically induce a more challenging optimization landscape and exacerbate model bias, so a short 4-step receding horizon empirically provides a practical sweet spot, and we replan frequently.
Contact-Directed Planning: Fig. 7 (a) visualizes how the model leverages its predictions for planning by showing the evolution of the objective function value in Eq. (11) for the hand’s target position. Early in the task (e.g., Step 30), the value map consistently encourages the hand to stay near a natural, energy-efficient default position. However, as the ball approaches and the plan solidifies (e.g., Step 60-70), a high-value region (yellow) emerges and sharpens, decisively guiding the robot’s hand toward the optimal contact point. This dynamic evolution of the value map showcases the model’s contact-directed reasoning, effectively forming an interpretable plan to achieve its objective.
@@ IV-E Real-World Validation
We deployed our method for real-time experiments on Unitree G1 with 25 Hz real-time planning and evaluated a batch of 1024 action trajectories with a planning horizon of 4 steps at each timestep. The desired base velocity $v$ was controlled by a human operator.
Our real-world deployments, shown in Fig. 4, included both single-task and multi-task models, both of which proved capable of completing their assigned tasks. The policy also demonstrated the ability to generalize to out-of-distribution (OOD) scenarios, such as blocking a previously unseen box. These experiments validate that our method can achieve agile and robust vision-based control. Crucially, the learned policy exhibits reactive, context-dependent behavior rather than overfitting to a single action pattern. For instance, in the “Support the Wall” task, the robot only braces its hands against the wall when actively disturbed and returns to a neutral stance once balance is recovered.
@@ V Conclusion
By integrating a scalable ego-centric visual world model with value-guided sampling-based MPC, we demonstrate that humanoid robots can efficiently and robustly learn agile, contact-rich behaviors from offline, demonstration-free data, advancing data-efficient, vision-based planning for real-world robotic interaction.
@@ ACKNOWLEDGMENT
This work was supported in part by NSF Grant CMMI-1944722 and by the Robotics and AI Institute. M. Ghaffari receives support from AFOSR MURI FA9550-23-1-0400. Z. Li was funded in part by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government via the Hong Kong Centre for Logistics Robotics. The authors would like to thank Jiaze Cai and Yen-Jen Wang for their help with experiments. We are also grateful to Bike Zhang, Fangchen Liu, Chaoyi Pan, Junfeng Long, and Yiyang Shao for their valuable discussions. This work was done during H. Liu’s visit to UC Berkeley.